OpenStudy (simplixity):

A rock is thrown straight down with an initial velocity of 15 m/s from the Verrazano Narrows Bridge in New York City. The roadway of this bridge is 70 m above water. Take upwards to be the positive direction. 1. Calculate the displacement at a time of 1.0 s. 2. Calculate the velocity at a time of 1.0 s.

OpenStudy (brainzonly):

The attempt at a solution: -5.6

OpenStudy (simplixity):

how did you get that? I have like 4 more problems just like this so I kinda wanna know how to get the answer

OpenStudy (brainzonly):

I've figured out the displacement by (10.5)(1)+(1/2)(9.8)(1.0)^2=-15.4 I don't know how to find velocity though.

OpenStudy (brainzonly):

wait! calculate the velocity at t = 1.0 s.

OpenStudy (brainzonly):

this is actually physics and math at the same time!!

OpenStudy (loser66):

do you know calculus?

OpenStudy (loser66):



OpenStudy (brainzonly):

yes I aced it why

OpenStudy (loser66):

You have the distance formula \(x = x_0+V_0t +(1/2) gt^2\) if you choose upward is + direction, then, the formula becomes \[X(t) = X_0-V_0t-(1/2) gt^2\]

OpenStudy (simplixity):

the -15.4 wasn't right

OpenStudy (loser66):

Now, At the beginning, \(X_0=70,V_0= 15m/s, g=9.8\) , and at t = 1.0 s, just plug in to get the answer for a

OpenStudy (brainzonly):

Since I was of no help, I am leaving. <<<<<<<<<<<

OpenStudy (loser66):

For velocity, if you know calculus, just take derivative of X w.r.t.t If you don't know calculus, this is the formula for velocity at time t \[V(t) = V_0+gt\] Minus sign to show that it is down ward.

OpenStudy (simplixity):

I got 50.095 for the displacement, is that right?

OpenStudy (brainzonly):

Well according to this source, no. But neither am I and this guy isn't a calculus professor..

OpenStudy (brainzonly):

However, these other 2 sources say yes you r correct

OpenStudy (simplixity):

@brainzonly haha i'm so confused

OpenStudy (simplixity):

I only have 1 more try at this problem, if I tell you the answer do you think you could help me figure out how they got it?

OpenStudy (brainzonly):

yes ill try

OpenStudy (brainzonly):

try 50.095 first

OpenStudy (simplixity):

I did, it was wrong

OpenStudy (brainzonly):

:(

OpenStudy (simplixity):

any other ideas? before I guess a number haha

OpenStudy (brainzonly):

nope. sorry.... @mathmale help!

OpenStudy (simplixity):

@Loser66 @brainzonly the answer for the displacement was -19.9, do you know how they got that?

OpenStudy (brainzonly):

give me a minute

OpenStudy (loser66):

\(x=x_0-V_0t-(1/2)gt^2\) replace all, you have \(x = 70 - 15*1-(1/2)9.8*(1)^2= 50.1\) but the question is about the displacement. That is this.